OpenStudy (anonymous):

the semi circles are drawn outside a triangle so that the sides of the triangle are the diameters of the semi circles. if the triangle is a right angled triangle and theses semi circles have areas 18, 32, and 50 determine the area of the triangle? @Mathematics

hero (hero):

1. Set A = (pir^2)/2 2. Find the radius 3. Multiply the radius by 2 to find the diameter 4. Repeat for the other two sides of the triangle 5. Verify that the sides satisfy pythagorean theorem 6. Use Heron's formula to find area of triangle

OpenStudy (anonymous):

can you show me how?

hero (hero):

Yes. If you have skype, it's easier. If not there are other means.

hero (hero):

I only mention skype because it has a screen sharing function

hero (hero):

Showing you on here would take forever

OpenStudy (anonymous):

i dont have one

hero (hero):

I wish you did

OpenStudy (anonymous):

show me the first step only

hero (hero):

Oops, you have to divide it by 2 before sqrt rooting it

hero (hero):

wait...

hero (hero):

A = (pir^2)/2 18 = (pir^2)/2 36 = pir^2 36/pi = r^2 sqrt{36/pi} = r

hero (hero):

So one of the sides of the triangle is 3.385

hero (hero):

i hope that makes sense

OpenStudy (anonymous):

not Really

hero (hero):

If you want to understand it, you must get skype because I can explain it to you as well as show you step by step how to do it. You will only need audio and chat. You don't even need a microphone to use skype and it is free to download and install. If you really want to understand this, you can get it and save yourself from all kinds of trouble. If you don't like it, you can just uninstall it.

hero (hero):

I've helped other students on here using skype and it is a really good tool